OpenStudy (anonymous):

intgration for sin4x*sin3x

OpenStudy (turingtest):

is this\[\int\sin(4x)\sin(3x)dx\]?

OpenStudy (anonymous):

yes

OpenStudy (turingtest):

use the identity\[\sin\alpha\sin\beta=\cos(\alpha-\beta)-\cos(\alpha+\beta)\]and it's a snap

OpenStudy (anonymous):

thank you very much

OpenStudy (turingtest):

welcome :)